# Inference-Time Rule Eraser: Fair Recognition via Distilling and Removing Biased Rules

### v3

@@ Abstract
Machine learning models often make predictions based on biased features such as gender, race, and other social attributes, posing significant fairness risks, especially in societal applications, such as hiring, banking, and criminal justice. Traditional approaches to addressing this issue involve retraining or fine-tuning neural networks with fairness-aware optimization objectives. However, these methods can be impractical due to significant computational resources, complex industrial tests, and the associated CO2 footprint. Additionally, regular users often fail to fine-tune models because they lack access to model parameters In this paper, we introduce the Inference-Time Rule Eraser (Eraser), a novel method designed to address fairness concerns by removing biased decision-making rules from deployed models during inference without altering model weights. We begin by establishing a theoretical foundation for modifying model outputs to eliminate biased rules through Bayesian analysis. Next, we present a specific implementation of Eraser that involves two stages: (1) distilling the biased rules from the deployed model into an additional patch model, and (2) removing these biased rules from the output of the deployed model during inference. Extensive experiments validate the effectiveness of our approach, showcasing its superior performance in addressing fairness concerns in AI systems.
@@ I Introduction
AArtificial intelligence (AI) systems have become increasingly prevalent in the real world and are widely deployed in many high-stakes applications such as face recognition and recruitment. However, these systems often inherit biases present in their training data, leading to unfair treatment based on protected attributes such as gender and race [1, 2, 3]. For example, the popular COMPAS algorithm for recidivism prediction exhibited bias against Black inmates and made unfair sentencing decisions [4]. Similarly, Microsoft’s face recognition system showed gender bias in classifying face attributes [5]. The prevalence of AI unfairness has surged in recent years, emphasizing the need to prioritize fairness alongside accuracy.
@@ II-A Fairness in Machine Learning
Fairness in machine learning (ML) has become a major research focus. Methods for measuring bias, such as Equalodds [9], have revealed that AI models can exhibit societal bias towards specific demographic groups, including gender and race [12, 13]. Strategies to promote fairness can be broadly classified into pre-processing, in-processing, and post-processing methods [14], which correspond to interventions on the training data, the training algorithm, and the trained model, respectively. Among them, in-processing has received the most research attention and has demonstrated the most effective debiasing performance, while post-processing methods have been the least studied.
Pre-processing methods are designed to modify the distribution of training data, making it difficult for the training algorithm to establish a statistical correlation between bias attributes and tasks [15, 16]. For example, study [15] suggests adjusting the distribution of training data to construct unbiased training data. Another study [16] proposes the generation of new samples to modify the distribution of training data through data augmentation. These strategies aim to decouple the target variable in the training data from the bias variable, thereby preventing the learning of bias in the training of the target task. In-processing methods focus on incorporating additional fairness-aware constraints during ML training [7, 6, 17, 18, 19, 20]. For instance, research [7] introduces adversarial fairness penalty terms, which minimize the learning of features related to bias variables during the learning process of the target task, thereby mitigating bias. Some studies [6, 17] propose fairness-aware contrastive learning loss that brings samples with different protected attributes closer in feature space, thereby reducing differential treatment by the model. Despite their effectiveness, these methods generally do not extend to debiasing deployed models without retraining.
@@
Our approach diverges from existing post-processing methods because we view the biases in the model as rules that can be removed. This novel approach allows us to systematically and flexibly eliminate bias in black-box models.
@@ II-B Rule Editing
Rule Editing [25] provides a lightweight approach to modifying the rules or knowledge embedded in pre-trained models, bypassing the need for full retraining. Existing methods for Rule Editing can be broadly classified into three categories: model fine-tuning, local parameter modification, and structural expansion.
Model fine-tuning is a straightforward approach to rule editing. It involves using data that complies with the desired rules to fine-tune the model. For instance, some studies [26] fine-tune pre-trained language models to correct inaccuracies in their knowledge, significantly reducing the computational costs compared to retraining. Other research [27] suggests further reducing costs by adjusting eigenvalues after performing singular value decomposition (SVD) on the backbone network. Research [25] proposes generating counterfactuals for erroneous rules in training data and using them for fine-tuning to correct the model’s rules. Local parameter modification [22, 28] focuses on adjusting specific model parameters related to the rules to be edited. These methods first identify the relevant parameters by comparing neuron activation differences between rule-aligned and rule-conflicting samples, then retrain these parameters. In addition to fine-tuning the identified local parameters, other research [29] proposes directly reversing the neural activation encoded by the located parameters without training. Structural expansion is another approach to address model editing. Research [30] leverages the properties of ReLU and proposes a theoretically guaranteed repair technique.
However, these methods require adjustments to the structure or parameter weights of white-box models, posing challenges for editing rules in black-box scenarios. In contrast, the Eraser method we propose only requires modifying the model’s output to indirectly edit its rules.
@@ III-A Problem Formulation
The problem of debiasing deployed models can be formally defined as follows: Given an already-trained deployed model, denoted as $\mathcal{M}$, which has been trained on a large dataset $\mathcal{D}_{L}$. This model $\mathcal{M}$ serves as a black-box machine learning service that takes an input $\mathbf{x}\in X$ and produces a probabilistic output $\mathcal{M}(\mathbf{x})$, where $\mathcal{M}(\mathbf{x})^{j}$ represents the probability of predicting $\mathbf{x}$ as the $j^{th}$ target class. The training dataset $\mathcal{D}_{L}$ and the model parameters are inaccessible. To perform debiasing on $\mathcal{M}$, we are provided with a small-scale dataset called the calibration set $\mathcal{D}_{S}=\{\mathbf{x}_{i},y_{i},b_{i}\}_{i=1:N},|\mathcal{D}_{S}|\ll|\mathcal{D}_{L}|$. Here $\mathbf{x}_{i}\in X_{S}$ denotes the $i^{th}$ input feature, $y_{i}\in Y$ denotes its target label, $b_{i}\in B$ denotes its bias label (e.g., gender), and $|\mathcal{D}_{S}|\ll|\mathcal{D}_{L}|$ represents the size of $\mathcal{D}_{S}$ is significantly smaller than the training set $\mathcal{D}_{L}$ of the deployed model. Under the constraint of not modifying the parameters of deployed model $\mathcal{M}$, the goal of debiasing deployed models is to ensure that the output is independent of the bias-related information in the input, while maintaining accuracy in predicting the target label.
@@ IV-A Overview
The Inference-Time Rule Eraser mitigates bias in deployed models by removing the biased rule $p(y|b)$ from the log space of the model’s output. Although the deployed model inherently includes $p(y|b)$, extracting this information directly from the deployed model presents challenges. To address this, we propose Rule Distillation Learning, inspired by the concept of Knowledge Distillation [31, 32]. This method comprises the following steps: (1) distilling the biased rule from the deployed model using a small calibration set based on concepts from causal inference, and (2) training a supplementary small model called a patch model to learn these rules.
@@ IV-B Rule Distillation learning
Rule Distillation.    To obtain $\mathcal{M}(X=\mathbf{x}^{b})$, we need to nullify the $\mathbf{x}^{y}$ feature in $\mathbf{x}$, resulting in an edited sample $\mathbf{x}^{(e)}$, and then input it into the model $\mathcal{M}$. However, given the difficulty of directly modifying the image $\mathbf{x}$ in pixel space, we adopt an indirect approach. We utilize $\mathbf{x}$ and contrastive samples to simulate sample editing at the representation layer of model $\mathcal{M}$, ensuring that only $\mathbf{x}^{b}$ is provided to the model $\mathcal{M}$. This approach is motivated by recent research indicating that features deep in a network tend to be linearized [34, 35]. As shown in Fig. 4(a), using a target class number $k$=2 as an example, there exists a semantic direction for modifying target features in the deep feature space (i.e., the model’s representation space $h(\cdot)$). We aim to simulate the result of feeding the edited sample $\mathbf{x}^{(e)}$ into the model $\mathcal{M}$. If the deep feature $h(\mathbf{x})$ of the original sample $\mathbf{x}$ moves along the aforementioned semantic direction towards a state away from any target class, this will result in maximum uncertainty regarding target information. The resulting deep feature can then be regarded as the deep feature representation $h(\mathbf{x}^{(e)})$ of $\mathbf{x}^{(e)}$.
@@ V-A Evaluation Setup
To answer the first question, we divide the test set into different groups based on target attributes and biased attributes, then assess accuracy for each group, and then report the average group accuracy and the worst group accuracy. To answer the second question, we examine whether the model predictions adhere to the fairness criterion Equalodds, which measures the changes in model predictions when the biased attributes shift. For example, in bignose and non-bignose recognition, EqualOdds measures gender bias as follows:
@@
Datasets.    We evaluate our method against the baselines across seven different datasets to ensure comprehensive coverage of various debiasing tasks. Seven of these datasets have been proposed in previous research and include both visual and tabular data. However, given the lack of datasets related to multi-class problems in industrial-scale data, we have additionally constructed ImageNet-B to test the debiasing performance.
CelebA [37] is a large-scale face image recognition dataset comprising 200,000 images, containing 40 attributes for each image, of which Gender is the bias attribute that should be prevented from being uesed for prediction. We chose Bignose, Blonde, and Attractive as target attributes for three recognition tasks to verify our gender debiasing effect, as they exhibit the highest Pearson correlation with gender. UTKface [38] is a dataset that contains around 20k facial images annotated with ethnicity and gender. It is used to validate ethnicity debiasing performance by constructing a biased training set with gender as the target attribute. We also used multi-class dataset C-MNIST [39] to evaluate debiasing. C-MNIST is a dataset of handwritten digits (0-9) with different background colors, where each digit has a strong correlation with the background color. Debiasing aims to eliminate the spurious dependency on background color information in the digit recognition task. In addition, we are interested in investigating the debiasing effect of our method when multiple biases coexist. The UrbanCar dataset [40] presents a unique opportunity as it encompasses two bias attributes: background (BG) and co-occurring object (Co-obj), with the target attribute being the type of car urban, country. This dataset was constructed by segmenting and subsequently recombining images from three distinct datasets: Stanford Cars [41] for car images, Places [42] for background images, and LVIS [43] for images of co-occurring objects. In addition to the aforementioned visual datasets, we also evaluated debiasing techniques on two structured/tabular datasets, namely LSAC [44] and COMPAS [45], both of which are extensively utilized in fairness literature. Specifically, we examined the performance of race debiasing in the context of law school admission prediction for LSAC, and recidivism prediction for COMPAS.
@@ V-B Quantitative Results
The debiasing performance observed in the multi-class datasets (C-MNIST and ImageNet-B), as presented in Tables I(e) and I(f), aligns with the results from the CelebA and UTKFace datasets. After applying our method, four diverse types of deployed models consistently show significant improvements in both accuracy and fairness. This indicates that our approach is not limited to binary tasks, which have been the primary focus of fairness research, but extends to more complex multi-class tasks as well. Moreover, the debiasing performance on ImageNet-B validates the suitability of our method for large-scale industrial datasets.
@@ V-E Analysis and discussion
TABLE VI: Results for the scenario using the LFW dataset as the calibration set of Eraser and the CelebA dataset as the test set.
Task Method Avg. Acc ($\uparrow$) Worst Acc ($\uparrow$) Model Bias ($\downarrow$)
BigNose Vanilla 69.48$\pm$0.2 36.45$\pm$0.6 23.26$\pm$0.3
Eraser 73.53$\pm$0.3 70.36$\pm$0.5 4.33$\pm$0.5
Blonde Vanilla 82.39$\pm$0.3 46.13$\pm$0.7 22.06$\pm$0.6
Eraser 89.37$\pm$0.6 84.41$\pm$0.8 6.72$\pm$0.4
Attrative Vanilla 76.85$\pm$0.4 66.20$\pm$0.7 20.92$\pm$0.4
Eraser 77.83$\pm$0.4 74.67$\pm$0.9 4.57$\pm$0.5
@@
Robustness to Calibration Dataset.    To evaluate the robustness of our proposed Eraser to the distribution of the calibration set, we conducted a qualitative analysis by training the patch model on one dataset and applying the patch model to another dataset. In other words, the calibration dataset and the test set come from different distributions. Specifically, we used the LFW [52] dataset as the calibration set to train the patch model, and the CelebA dataset as the test set. The results of deployed models before and after integration with the Eraser are shown in Table VI.
[…]
@@ VI Conclusion
This paper introduced the Inference-Time Rule Eraser (Eraser), a novel approach to fairness in AI systems. Unlike existing methods, Eraser does not require access to or modification of model weights. It operates by removing biased rules at inference time, which is achieved by modifying the model’s output. We also proposed a specific implementation, which distills biased rules from the deployed model into an additional patched model and removes them during inference time. Our experiments demonstrated the effectiveness and superior performance of Eraser.

### v4

@@ Abstract
Machine learning models often make predictions based on biased features such as gender, race, and other social attributes, posing significant fairness risks, especially in societal applications, such as hiring, banking, and criminal justice. Traditional approaches to addressing this issue involve retraining or fine-tuning neural networks with fairness-aware optimization objectives. However, these methods can be impractical due to significant computational resources, complex industrial tests, and the associated CO2 footprint. Additionally, regular users often fail to fine-tune models because they lack access to model parameters. In this paper, we introduce the Inference-Time Rule Eraser (Eraser), a novel method designed to address fairness concerns by removing biased decision-making rules from deployed models during inference without altering model weights. We begin by establishing a theoretical foundation for modifying model outputs to eliminate biased rules through Bayesian analysis. Next, we present a specific implementation of Eraser that involves two stages: (1) distilling the biased rules from the deployed model into an additional patch model, and (2) removing these biased rules from the output of the deployed model during inference. Extensive experiments validate the effectiveness of our approach, showcasing its superior performance in addressing fairness concerns in AI systems.
@@ I Introduction
AArtificial intelligence (AI) systems have become increasingly prevalent in the real world and are widely deployed in many high-stakes applications such as face recognition and recruitment. However, these systems often inherit biases present in their training data, leading to unfair treatment based on protected attributes such as gender and race [1, 2]. For example, the popular COMPAS algorithm for recidivism prediction exhibited bias against Black inmates and made unfair sentencing decisions [3]. Similarly, Microsoft’s face recognition system showed gender bias in classifying face attributes [4]. The prevalence of AI unfairness has surged in recent years, emphasizing the need to prioritize fairness alongside accuracy.
@@ II-A Fairness in Machine Learning
The problem of model bias in deep learning has received increasing attention in recent years [11, 12, 13, 14, 15, 16, 17]. Methods for measuring bias, such as Equalodds [8], have revealed that AI models can exhibit societal bias towards specific demographic groups, including gender and race [18, 19]. Strategies to promote fairness can be broadly classified into pre-processing, in-processing, and post-processing methods [20], which correspond to interventions on the training data, the training algorithm, and the trained model, respectively. Among them, in-processing has received the most research attention and has demonstrated the most effective debiasing performance, while post-processing methods have been the least studied.
Pre-processing methods are designed to modify the distribution of training data, making it difficult for the training algorithm to establish a statistical correlation between bias attributes and tasks [21, 22]. For example, study [21] suggests adjusting the distribution of training data to construct unbiased training data. Another study [22] proposes the generation of new samples to modify the distribution of training data through data augmentation. These strategies aim to decouple the target variable in the training data from the bias variable, thereby preventing the learning of bias in the training of the target task. In-processing methods focus on incorporating additional fairness-aware constraints during ML training [6, 5, 23]. For instance, research [6] introduces adversarial fairness penalty terms, which minimize the learning of features related to bias variables during the learning process of the target task, thereby mitigating bias. Some studies [5, 23] propose fairness-aware contrastive learning loss that brings samples with different protected attributes closer in feature space, thereby reducing differential treatment by the model. Despite their effectiveness, these methods generally do not extend to debiasing deployed models without retraining.
@@
Unlike existing methods, we investigate a debiasing approach that relies solely on model output, without requiring access to model parameters or any labels of test samples.
@@ II-B Rule Editing
Rule Editing [28] provides a lightweight approach to modifying the rules or knowledge embedded in pre-trained models, bypassing the need for full retraining.
Model fine-tuning is a straightforward approach to rule editing. It involves using data that complies with the desired rules to fine-tune the model. For instance, some studies [29] fine-tune pre-trained language models to correct inaccuracies in their knowledge, significantly reducing the computational costs compared to retraining. Other research [30] suggests further reducing costs by adjusting eigenvalues after performing singular value decomposition (SVD) on the backbone network. Research [28] proposes generating counterfactuals for erroneous rules in training data and using them for fine-tuning to correct the model’s rules. Local parameter modification [25, 31] focuses on adjusting specific model parameters related to the rules to be edited. In addition to fine-tuning the identified local parameters, other research [32] proposes directly reversing the neural activation encoded by the located parameters without training. Structural expansion is another approach to address model editing. Research [33] leverages the properties of ReLU and proposes a theoretically guaranteed repair technique.
However, these methods require adjustments to the structure or parameter weights of white-box models, posing challenges for editing rules in black-box scenarios.
@@ III-A Problem Formulation
The problem of debiasing deployed models can be formally defined as follows: Given an already-trained deployed model, denoted as $\mathcal{M}$, which has been trained on a large dataset $\mathcal{D}_{L}$. This model $\mathcal{M}$ serves as a black-box machine learning service that takes an input $\mathbf{x}\in X$ and produces a probabilistic output $\mathcal{M}(\mathbf{x})$, where $\mathcal{M}(\mathbf{x})^{j}$ represents the probability of predicting $\mathbf{x}$ as the $j^{th}$ target class. The training dataset $\mathcal{D}_{L}$ and the model parameters are inaccessible. To perform debiasing on $\mathcal{M}$, we are provided with a small-scale dataset called the calibration set $\mathcal{D}_{S}=\{\mathbf{x}_{i},y_{i},b_{i}\}_{i=1:N},|\mathcal{D}_{S}|\ll|\mathcal{D}_{L}|$. Here $\mathbf{x}_{i}\in X_{S}$ denotes the $i^{th}$ input feature, $y_{i}\in Y$ denotes its target label, $b_{i}\in B$ denotes its bias label (e.g., gender), and $|\mathcal{D}_{S}|\ll|\mathcal{D}_{L}|$ represents the size of $\mathcal{D}_{S}$ is significantly smaller than the training set $\mathcal{D}_{L}$ of the deployed model. Subject to the constraint of not altering the parameters of the deployed model $\mathcal{M}$, the objective of debiasing is to ensure that the model’s output remains independent of bias-related information in the input, while preserving accuracy in predicting the target label.
@@ IV-A Overview
The Inference-Time Rule Eraser mitigates bias in deployed models by removing the biased rule $p(y|b)$ from the log space of the model’s output. Although the deployed model inherently includes $p(y|b)$, extracting this information directly from the deployed model presents challenges. To address this, we propose Rule Distillation Learning, inspired by the concept of Knowledge Distillation. This method comprises the following steps: (1) distilling the biased rule from the deployed model using a small calibration set based on concepts from causal inference, and (2) training a supplementary small model called a patch model to learn these rules.
@@ IV-B Rule Distillation learning
Rule Distillation.    To obtain $\mathcal{M}(X=\mathbf{x}^{b})$, we need to nullify the $\mathbf{x}^{y}$ feature in $\mathbf{x}$, resulting in an edited sample $\mathbf{x}^{(e)}$, and then input it into the model $\mathcal{M}$. However, given the difficulty of directly modifying the image $\mathbf{x}$ in pixel space, we adopt an indirect approach. We utilize $\mathbf{x}$ and contrastive samples to simulate sample editing at the representation layer of model $\mathcal{M}$, ensuring that only $\mathbf{x}^{b}$ is provided to the model $\mathcal{M}$. This approach is motivated by recent research indicating that features deep in a network tend to be linearized [35, 36]. As shown in Fig. 4(a), using a target class number $k$=2 as an example, there exists a semantic direction for modifying target features in the deep feature space (i.e., the model’s representation space $h(\cdot)$). We aim to simulate the result of feeding the edited sample $\mathbf{x}^{(e)}$ into the model $\mathcal{M}$. If the deep feature $h(\mathbf{x})$ of the original sample $\mathbf{x}$ moves along the aforementioned semantic direction towards a state away from any target class, this will result in maximum uncertainty regarding target information. The resulting deep feature can then be regarded as the feature representation $h(\mathbf{x}^{(e)})$ of $\mathbf{x}^{(e)}$.
@@ V-A Evaluation Setup
To answer the first question, we divide the test set into different groups based on target attributes and bias attributes, then assess accuracy for each group, and then report the average group accuracy and the worst group accuracy. To answer the second question, we examine whether the model predictions adhere to the fairness criterion Equalodds, which measures the changes in model predictions when the bias attributes shift. For example, in bignose and non-bignose recognition, EqualOdds measures gender bias as follows:
@@
Datasets.    We evaluate our method against the baselines across seven different datasets to ensure comprehensive coverage of various debiasing tasks. However, given the lack of datasets related to multi-class problems in industrial-scale data, we have additionally constructed ImageNet-B to test the debiasing performance.
CelebA [38] is a large-scale face image recognition dataset comprising 200,000 images, containing 40 attributes for each image, of which Gender is the bias attribute that should be prevented from being uesed for prediction. We chose Bignose, Blonde, and Attractive as target attributes for three recognition tasks to verify our gender debiasing effect, as they exhibit the highest Pearson correlation with gender. UTKface [39] is a dataset that contains around 20k facial images annotated with ethnicity and gender. It is used to validate ethnicity debiasing performance by constructing a biased training set with gender as the target attribute. We also used multi-class dataset C-MNIST [40] to evaluate debiasing. C-MNIST is a dataset of handwritten digits (0-9) with different background colors, where each digit has a strong correlation with the background color. Debiasing aims to eliminate the spurious dependency on background color information in the digit recognition task. In addition, we are interested in investigating the debiasing effect of our method when multiple biases coexist. The UrbanCar dataset [41] presents a unique opportunity as it encompasses two bias attributes: background (BG) and co-occurring object (Co-obj), with the target attribute being the type of car urban, country. In addition to the aforementioned visual datasets, we also evaluated debiasing techniques on two structured/tabular datasets, namely LSAC [42] and COMPAS [43], both of which are extensively utilized in fairness literature. Specifically, we examined the performance of race debiasing in the context of law school admission prediction for LSAC, and recidivism prediction for COMPAS.
@@ V-B Quantitative Results
The debiasing performance observed in the multi-class datasets (C-MNIST and ImageNet-B), as presented in Tables I(e) and I(f), aligns with the results from the CelebA and UTKFace datasets. After applying our method, four diverse types of deployed models consistently show significant improvements in both accuracy and fairness.
@@ VI Conclusion
This paper introduced the Inference-Time Rule Eraser (Eraser), a novel approach to enhancing fairness in AI systems without requiring access to or modification of model weights. Eraser functions by eliminating biased rules during inference through output modification. We also presented a specific implementation that distills biased rules from the deployed model into a supplementary patched model, which removes them during inference. Experimental results validate the effectiveness and superiority of Eraser in mitigating bias.